# Search for neutrino emission from blazar $\gamma$-ray flares accounting for possible neutrino time delays

### v1

@@ 1 Introduction
In the 2010s, the IceCube neutrino observatory discovered astrophysical high-energy ($E_{\nu}\gtrsim 1$ TeV) neutrinos [1]. Around $\sim 5-30$% of their flux was later attributed to the Milky Way galaxy [2, 3, 4]. Aside from that, besides the evidence of an individual association of the multi-wavelength flare in blazar TXS 0506+056 with a $\gtrsim 200$ TeV neutrino [5] and evidence of 0.31000.3-100 TeV neutrino emission from active galaxies NGC 1068 [6], NGC 3079 and NGC 4151 [7], the main contributors to the observed neutrino flux remain unknown (for a recent review, see, e.g., [8, 9, 10]).
Active galactic nuclei (AGNs) and their subset with relativistic jets pointing to the Earth at a small angle, blazars, were proposed as neutrino sources almost half a century ago [11]. Still, it was not until recently when some evidence of their detectable contribution to the observed high-energy neutrino flux has been found [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Yet, the (in)significance of the blazar-neutrino association depends on neutrino datasets, blazar catalogues and test statistics used in searches for correlations between AGNs and neutrinos [24, 25].
Previous studies, looking for associations between blazars and IceCube high-energy neutrinos, used as proxies either time-averaged fluxes of sources in various electromagnetic bands (from radio to $\gamma$-ray) or their light curves with corresponding values close to the time of the neutrino arrival. However, from the theoretical point of view, the temporal coincidence of the neutrino arrival and the major electromagnetic flare (here, we focus on $\gamma$-ray flares; note that radio flares tend to lag behind $\gamma$-ray flares [26] and have longer duration, so we do not discuss them here) might not be warranted for at least two reasons: (i) slow proton acceleration, and/or (ii) slow proton energy losses.
@@
Whether it is the acceleration timescale $t^{\prime}_{\mathrm{acc}}$ or the $p\gamma$ energy-loss timescale $t^{\prime}_{p\gamma}$ which causes the possible time lag $t^{\prime}_{\mathrm{delay}}$ of the neutrino arrival depends on the conditions in the source, such as the acceleration mechanism, magnetic field, and target photon fields, and it is the slowest process which will determine the neutrino time lag, i.e. $t^{\prime}_{\mathrm{delay}}\sim\max\{t^{\prime}_{\mathrm{acc}},\,t^{\prime}_{p\gamma}\}$. Depending on the model parameters, $\gtrsim 100$ TeV neutrinos can arrive months to years later than the corresponding electromagnetic flare peak [27]. A study of a possible association of a few lagging neutrinos with another class of sources, namely tidal disruption events, was recently performed by [28].
If one argues that the proton acceleration timescale is not substantially larger than the electron acceleration timescale, and the $p\gamma$ collisions happen far from the threshold and efficiently, then protons can be accelerated to energies as high as $\sim 100$ PeV and produce neutrinos of corresponding energies (with a negligible time delay with respect to the electromagnetic flare), since the peak energy of the neutrino spectral-energy distribution closely follows the maximum proton energy [27, 29]. However, only one neutrino with energy $E_{\nu}\gtrsim$100 PeV has been discovered so far, namely KM3-230213A [30]. Thus, some of those $\gtrsim 100$ TeV astrophysical neutrinos detected by the IceCube observatory [31], if they are produced in a causal connection with major flares in blazars, can be expected to come with some time delay with respect to the flare peaks, either due to the large proton acceleration or $p\gamma$ energy-loss timescale.
The publicly available Fermi-LAT light-curve repository (LCR) [32] with regular-binned light curves of all significantly variable Fermi-LAT-detected AGNs spanning more than 15 years allows us to search for association of major $\gamma$-ray flares with lagging behind them IceCube high-energy neutrinos published by the IceCube collaboration in the first alert catalogue IceCat-1 [31, 33]. In this paper, we make the first attempt at this search. In Sect. 2, we present the datasets and catalogues used in our work; in Sect. 3, we present the methods used for searching for associations of lagging neutrinos with major $\gamma$-ray flares in AGNs; in Sect. 4, the results are shown. We discuss implications of our findings and conclude in Sect. 5.
@@ 2.1 Neutrino alerts from the IceCat-1 catalogue
We use the IceCat-1 catalogue of high-energy neutrino alerts published by the IceCube collaboration [31, 33]. In the following, we use the terms “alert” and “neutrino” interchangeably. We remove eight entries for which there is a flag indicating a coinciding signal in the IceTop detector of cosmic rays, implying a high chance of those alerts being background. This leaves us with $N_{\nu}=348$ neutrino alerts. For each alert $\nu$, the following information of interest is available:
$x_{\nu}=\{s_{\nu},t_{\nu},E_{\nu},\alpha_{\nu},\delta_{\nu},\vec{\epsilon}_{\nu}\},$ (2.1)
where $s_{\nu}$ is the neutrino signalness reported by the IceCube collaboration in the IceCat-1 catalogue which varies from $0.126$ to $0.997$ with a median value of $0.410$ and represents the probability of the neutrino $\nu$ being of astrophysical origin assuming an $E^{-2.19}_{\nu}$ astrophysical neutrino power-law flux (although this assumption does not hold, e.g., in cases when neutrinos are produced in $p\gamma$ interactions near the threshold, see, e.g., [34, 35, 36]), $t_{\nu}$ is the moment of the neutrino detection, $E_{\nu}$ is the estimated muon energy, which is a good proxy of the neutrino energy (see Fig. 7 in Ref. [31]), $\alpha_{\nu}$ is the right ascension (RA), and $\delta_{\nu}$ is the declination of the most probable sky coordinate of the neutrino arrival, while the vector $\vec{\epsilon}_{\nu}=\{\alpha_{\nu}^{+},\,\alpha_{\nu}^{-},\,\delta_{\nu}^{+},\,\delta_{\nu}^{-}\}$ incorporates information about the uncertainty of the reconstructed direction of the neutrino arrival, where $(\alpha_{\nu}^{+}+\alpha_{\nu}^{-})$ defines the width, and $(\delta_{\nu}^{+}+\delta_{\nu}^{-})$ defines the length of the minimum rectangle encapsulating the neutrino uncertainty arrival contours (see Fig. 1 in Ref. [25]).
@@ 2.1.1 Set of mock neutrino datasets
$m_{\nu}^{f}=\{s_{\nu},t_{\nu}^{f},E_{\nu},\alpha_{\nu}^{f},\delta_{\nu},\vec{\epsilon}_{\nu}\}.$ (2.2)
Since the IceCube detector is located at the South Pole, its sensitivity to astrophysical neutrinos on a long timescale depends only on the energy and declination. Thus, following, e.g. [12], by generating a random uniformly distributed number between $0^{\circ}$ and $360^{\circ}$ for each of the values of $\alpha_{\nu}^{f}$, we obtain a set of mock neutrino datasets reflecting a situation of neutrinos registered from random sky directions (losing a possible connection with a potential class of sources) but keeping the dependence of the IceCube sensitivity on the declination unchanged. For correlation searches involving the time of the neutrino arrival, we generate random values of $t_{\nu}^{f}$ from a uniform distribution $U[t_{\nu}^{\mathrm{min}},t_{\nu}^{\mathrm{max}}]$, where $t_{\nu}^{\mathrm{min}}=55695.064$ (MJD) is the arrival time of the first neutrino in the IceCat-1, and $t_{\nu}^{\mathrm{max}}=60231.917$ (MJD) is the time of the last alert in IceCat-1. Thus, the total time span of IceCat-1 is $T_{\mathrm{\textit{IceCube}}}\approx 4537$ d $\approx 12.4$ yr. Using the Kolmogorov-Smirnov test, we verified, that the original set of RAs of IceCube neutrinos $\{\alpha_{\nu}\}$ and their arrival times $\{t_{\nu}\}$ cannot be distinguished from a set drawn from the corresponding uniform distribution.
@@ 2.2.1 MOJAVE-based source list
Since the aim of the paper is to search for blazar-neutrino associations with account for a possible neutrino lagging, which we assume to be universal in the jet frame, to transform the anticipated neutrino arrival delay from the jet frame into the observer frame, the measurements of the blazar’s Doppler factor $D$ and redshift $z$ are required, since $t_{\mathrm{\,delay}}=t^{\prime}_{\mathrm{delay}}(1+z)/D$. As a starting point, we use the sample of $447$ AGNs compiled in [37], which consists of radio-bright AGNs monitored by the MOJAVE program. All sources in the selected list are located at declinations $30^{\circ}$ and have a minimum 15 GHz correlated flux density $\gtrsim 50$ mJy. All these sources are also present in the Radio Fundamental Catalogue (RFC), version 2025a [38]. We utilize Table 4 in [37], where the Doppler factors of the blazars are reported, obtained under the assumption that all sources have observed brightness temperature $\propto DT^{\prime}_{b}/(1+z)$, and the intrinsic brightness temperatures $T^{\prime}_{b}$ are determined according to [39].
@@
As the next step, we use the 4FGL-DR4 catalogue of Fermi-LAT sources [40] available at the mission website [^3] https://fermi.gsfc.nasa.gov/ssc/data/access/lat/14yr_catalog/gll_psc_v35.fit as the file gll_psc_v35.fit. We cross-match 4FGL sources with those presented in the catalogue of Ref. [37].
@@
We use the 4LAC catalogue [41] to obtain the redshift values for the sources in our list of blazars. For those objects which do not have a finite redshift value in the 4LAC catalogue, we use the astroquery.ned module to query the redshift from the NASA/IPAC Extragalactic Database [42].
@@
Finally, we use the Fermi-LAT LCR [32] and download monthly-binned photon flux light curves. We leave only those objects which are present in the LCR, i.e. those with significant variability. The final list of 4FGL AGNs with known redshifts, $\gamma$-ray light curves and Doppler factors reported in [37] consists of $N_{b}=294$ sources. Among them, according to the classification in the 4FGL catalogue, there are 184 flat-spectrum radio quasars (FSRQs), 93 BL Lacertae-like objects (BLLs), 6 radio galaxies (RDGs), 4 blazars of uncertain type (BCUs), 4 Seyfert galaxies of type 1 (NLSY1s), 2 compact steep-spectrum radio sources (CSSs), and 1 steep-spectrum radio quasar (SSRQ). Since most of the sources are blazars, we use the terms “source”, “AGN”, and “blazar” interchangeably throughout the paper. For blazars for which only lower limits on Doppler factors were reported in [37] due to unknown redshift, we multiply the lower limits on $D$ by $(1+z)$ with the values of $z$ we found in the NASA/IPAC Extragalactic Database [^4] This is because the observed brightness temperature $\propto DT^{\prime}_{b}/(1+z)$ was assumed to correspond to $z=0$ for sources with unknown redshifts in [37], hence the lower limit on $D$ was reported.. The obtained blazar list is denoted as H21+.
@@ 2.2.2 CGRaBS-based source list
Doppler factors in [37] are obtained under the assumption that the apparent brightness temperature at $15$ GHz depends only on the blazar jet Doppler factor $D$ and the intrinsic (jet-frame) brightness temperature, which for all blazars in the sample is equal to $4\times 10^{10}$ K, as found according to [39]. This approach can give only approximate values of the Doppler factor because it does not allow for intrinsic source-to-source variations and, besides, may result in incorrect estimations of $D$ when radio images of the jet and the core of the blazar are superimposed [43].
Thus, it is of interest to search for blazar-neutrino associations using a sample of sources with their Doppler factors obtained with a different method. Recently, Ref. [34] modelled time-averaged spectral-energy distributions (SEDs) for a sample of 324 blazars, obtaining a fit to the multi-wavelength SEDs with a leptonic and a leptohadronic model, determining values of the Doppler factor $D$ (among other parameters) for each source. The source sample of Ref. [34] is based on the flux-limited ($F_{8.4\mathrm{\,GHz}}>65$ mJy) CGRaBS catalogue of radio-loud AGNs [44] which was cross-matched by [45] with sources from various Fermi-LAT catalogues.
We cross-match sources from the list of 324 blazars [34] [^5] Their Doppler factors and redshifts available at https://github.com/xrod/lephad-blazars/blob/main/model_parameters.csv. with the Fermi-LAT LCR [32]. This results in $N_{b}=295$ sources with known $D$, $z$ and light curves in the LCR. Among them, according to the Fermi-LAT classification [40], there are $218$ FSRQs, $72$ BLLs, $2$ BCUs, and $3$ NLSY1s. We denote the compiled source list as R24+.

### v2

@@ Abstract
keywords:
active galactic nuclei, neutrino astronomy
@@ 1 Introduction
In the 2010s, the IceCube neutrino observatory discovered astrophysical high-energy ($E_{\nu}\gtrsim 1$ TeV) neutrinos (Aartsen and others, 2013). Around $\sim 5-30$% of their flux was later attributed to the Milky Way galaxy (Abbasi and others, 2023b; Kovalev et al., 2022; Allakhverdyan and others, 2025). Aside from that, besides the evidence of an individual association of the multi-wavelength flare in blazar TXS 0506+056 with a $\gtrsim 200$ TeV neutrino (Aartsen and others, 2018) and evidence of 0.31000.3-100 TeV neutrino emission from active galaxies NGC 1068 (Abbasi and others, 2022), NGC 3079 and NGC 4151 (Neronov et al., 2024), the main contributors to the observed neutrino flux remain unknown (for a recent review, see, e.g., Troitsky, 2021, 2024; Kurahashi et al., 2022).
Active galactic nuclei (AGNs) and their subset with relativistic jets pointing to the Earth at a small angle, blazars, were proposed as neutrino sources almost half a century ago (Berezinsky and Ginzburg, 1981). Still, it was not until recently when some evidence of their detectable contribution to the observed high-energy neutrino flux has been found (Plavin et al., 2020, 2021, 2023, 2024, 2025; Hovatta and others, 2021; Giommi et al., 2020; Kouch and others, 2024; Buson and others, 2022; Buson et al., 2022; Suray and Troitsky, 2023; Albert and others, 2024). Yet, the (in)significance of the blazar-neutrino association depends on neutrino datasets, blazar catalogues and test statistics used in searches for correlations between AGNs and neutrinos (Bellenghi et al., 2023; Abbasi and others, 2023c).
Previous studies, looking for associations between blazars and IceCube high-energy neutrinos, used as proxies either time-averaged fluxes of sources in various electromagnetic bands (from radio to $\gamma$-ray) or their light curves with corresponding values close to the time of the neutrino arrival. However, from the theoretical point of view, the temporal coincidence of the neutrino arrival and the major electromagnetic flare (here, we focus on $\gamma$-ray flares; note that radio flares tend to lag behind $\gamma$-ray flares (Kramarenko et al., 2021) and have longer duration, so we do not discuss them here) might not be warranted for at least two reasons: (i) slow proton acceleration, and/or (ii) slow proton energy losses.
@@
Whether it is the acceleration timescale $t^{\prime}_{\mathrm{acc}}$ or the $p\gamma$ energy-loss timescale $t^{\prime}_{p\gamma}$ which causes the possible time lag $t^{\prime}_{\mathrm{delay}}$ of the neutrino arrival depends on the conditions in the source, such as the acceleration mechanism, magnetic field, and target photon fields, and it is the slowest process which will determine the neutrino time lag, i.e. $t^{\prime}_{\mathrm{delay}}\sim\max\{t^{\prime}_{\mathrm{acc}},\,t^{\prime}_{p\gamma}\}$. Depending on the model parameters, $\gtrsim 100$ TeV neutrinos can arrive months to years later than the corresponding electromagnetic flare peak (Podlesnyi and Oikonomou, 2025). A study of a possible association of a few lagging neutrinos with another class of sources, namely tidal disruption events, was recently performed by Lu et al. (2025).
If one argues that the proton acceleration timescale is not substantially larger than the electron acceleration timescale, and the $p\gamma$ collisions happen far from the threshold and efficiently, then protons can be accelerated to energies as high as $\sim 100$ PeV and produce neutrinos of corresponding energies (with a negligible time delay with respect to the electromagnetic flare), since the peak energy of the neutrino spectral-energy distribution closely follows the maximum proton energy (Podlesnyi and Oikonomou, 2025; Rodrigues et al., 2025). However, only one neutrino with energy $E_{\nu}\gtrsim$100 PeV has been discovered so far, namely KM3-230213A (KM3NeT Collaboration, 2025). Thus, some of those $\gtrsim 100$ TeV astrophysical neutrinos detected by the IceCube observatory (Abbasi and others, 2023a), if they are produced in a causal connection with major flares in blazars, can be expected to come with some time delay with respect to the flare peaks, either due to the large proton acceleration or $p\gamma$ energy-loss timescale.
The publicly available Fermi-LAT light-curve repository (LCR) (Abdollahi and others, 2023) with regular-binned light curves of all significantly variable Fermi-LAT-detected AGNs spanning more than 15 years allows us to search for association of major $\gamma$-ray flares with lagging behind them IceCube high-energy neutrinos published by the IceCube collaboration in the first alert catalogue IceCat-1 (Abbasi and others, 2023a; IceCube Collaboration, 2023). In this paper, we make the first attempt at this search. In Sect. 2, we present the datasets and catalogues used in our work; in Sect. 3, we present the methods used for searching for associations of lagging neutrinos with major $\gamma$-ray flares in AGNs; in Sect. 4, the results are shown. We discuss implications of our findings and conclude in Sect. 5.
@@ 2.1 Neutrino alerts from the IceCat-1 catalogue
We use the IceCat-1 catalogue of high-energy neutrino alerts published by the IceCube collaboration (Abbasi and others, 2023a; IceCube Collaboration, 2023). In the following, we use the terms “alert” and “neutrino” interchangeably. We remove eight entries for which there is a flag indicating a coinciding signal in the IceTop detector of cosmic rays, implying a high chance of those alerts being background. This leaves us with $N_{\nu}=348$ neutrino alerts. For each alert $\nu$, the following information of interest is available:
$x_{\nu}=\{s_{\nu},t_{\nu},E_{\nu},\alpha_{\nu},\delta_{\nu},\vec{\epsilon}_{\nu}\},$ (1)
where $s_{\nu}$ is the neutrino signalness reported by the IceCube collaboration in the IceCat-1 catalogue which varies from $0.126$ to $0.997$ with a median value of $0.410$ and represents the probability of the neutrino $\nu$ being of astrophysical origin assuming an $E^{-2.19}_{\nu}$ astrophysical neutrino power-law flux (although this assumption does not hold, e.g., in cases when neutrinos are produced in $p\gamma$ interactions near the threshold, see, e.g., Rodrigues et al., 2024a, b; Kuhlmann and Capel, 2025), $t_{\nu}$ is the moment of the neutrino detection, $E_{\nu}$ is the estimated muon energy, which is a good proxy of the neutrino energy (see Fig. 7 by Abbasi and others, 2023a), $\alpha_{\nu}$ is the right ascension (RA), and $\delta_{\nu}$ is the declination of the most probable sky coordinate of the neutrino arrival, while the vector $\vec{\epsilon}_{\nu}=\{\alpha_{\nu}^{+},\,\alpha_{\nu}^{-},\,\delta_{\nu}^{+},\,\delta_{\nu}^{-}\}$ incorporates information about the uncertainty of the reconstructed direction of the neutrino arrival, where $(\alpha_{\nu}^{+}+\alpha_{\nu}^{-})$ defines the width, and $(\delta_{\nu}^{+}+\delta_{\nu}^{-})$ defines the length of the minimum rectangle encapsulating the neutrino uncertainty arrival contours (see Fig. 1 by Abbasi and others, 2023c).
@@ 2.1.1 Set of mock neutrino datasets
$m_{\nu}^{f}=\{s_{\nu},t_{\nu}^{f},E_{\nu},\alpha_{\nu}^{f},\delta_{\nu},\vec{\epsilon}_{\nu}\}.$ (2)
Since the IceCube detector is located at the South Pole, its sensitivity to astrophysical neutrinos on a long timescale depends only on the energy and declination. Thus, following Plavin et al. (2020), by generating a random uniformly distributed number between $0^{\circ}$ and $360^{\circ}$ for each of the values of $\alpha_{\nu}^{f}$, we obtain a set of mock neutrino datasets reflecting a situation of neutrinos registered from random sky directions (losing a possible connection with a potential class of sources) but keeping the dependence of the IceCube sensitivity on the declination unchanged. For correlation searches involving the time of the neutrino arrival, we generate random values of $t_{\nu}^{f}$ from a uniform distribution $U[t_{\nu}^{\mathrm{min}},t_{\nu}^{\mathrm{max}}]$, where $t_{\nu}^{\mathrm{min}}=55695.064$ (MJD) is the arrival time of the first neutrino in the IceCat-1, and $t_{\nu}^{\mathrm{max}}=60231.917$ (MJD) is the time of the last alert in IceCat-1. Thus, the total time span of IceCat-1 is $T_{\mathrm{\textit{IceCube}}}\approx 4537$ d $\approx 12.4$ yr. Using the Kolmogorov-Smirnov test, we verified, that the original set of RAs of IceCube neutrinos $\{\alpha_{\nu}\}$ and their arrival times $\{t_{\nu}\}$ cannot be distinguished from a set drawn from the corresponding uniform distribution.
@@ 2.2.1 MOJAVE-based source list
Since the aim of the paper is to search for blazar-neutrino associations with account for a possible neutrino lagging, which we assume to be universal in the jet frame, to transform the anticipated neutrino arrival delay from the jet frame into the observer frame, the measurements of the blazar’s Doppler factor $D$ and redshift $z$ are required, since $t_{\mathrm{\,delay}}=t^{\prime}_{\mathrm{delay}}(1+z)/D$. As a starting point, we use the sample of $447$ AGNs compiled by Homan et al. (2021), which consists of radio-bright AGNs monitored by the MOJAVE program. All sources in the selected list are located at declinations $30^{\circ}$ and have a minimum 15 GHz correlated flux density $\gtrsim 50$ mJy. All these sources are also present in the Radio Fundamental Catalogue (RFC), version 2025a (Petrov and Kovalev, 2025). We utilize Table 4 in the work of Homan et al. (2021), where the Doppler factors of the blazars are reported, obtained under the assumption that all sources have observed brightness temperature $\propto DT^{\prime}_{b}/(1+z)$, and the intrinsic brightness temperatures $T^{\prime}_{b}$ are determined according to Kovalev and others (2005).
@@
As the next step, we use the 4FGL-DR4 catalogue of Fermi-LAT sources (Abdollahi and others, 2020) available at the mission website [^2] https://fermi.gsfc.nasa.gov/ssc/data/access/lat/14yr_catalog/gll_psc_v35.fit as the file gll_psc_v35.fit. We cross-match 4FGL sources with those presented in the catalogue of Homan et al. (2021).
@@
We use the 4LAC catalogue (Ajello and others, 2020) to obtain the redshift values for the sources in our list of blazars. For those objects which do not have a finite redshift value in the 4LAC catalogue, we use the astroquery.ned module to query the redshift from the NASA/IPAC Extragalactic Database (NASA/IPAC Extragalactic Database, 2019).
@@
Finally, we use the Fermi-LAT LCR (Abdollahi and others, 2023) and download monthly-binned photon flux light curves. We leave only those objects which are present in the LCR, i.e. those with significant variability. The final list of 4FGL AGNs with known redshifts, $\gamma$-ray light curves and Doppler factors reported by Homan et al. (2021) consists of $N_{b}=294$ sources. Among them, according to the classification in the 4FGL catalogue, there are 184 flat-spectrum radio quasars (FSRQs), 93 BL Lacertae-like objects (BLLs), 6 radio galaxies (RDGs), 4 blazars of uncertain type (BCUs), 4 Seyfert galaxies of type 1 (NLSY1s), 2 compact steep-spectrum radio sources (CSSs), and 1 steep-spectrum radio quasar (SSRQ). Since most of the sources are blazars, we use the terms “source”, “AGN”, and “blazar” interchangeably throughout the paper. For blazars for which only lower limits on Doppler factors were reported by Homan et al. (2021) due to unknown redshift, we multiply the lower limits on $D$ by $(1+z)$ with the values of $z$ we found in the NASA/IPAC Extragalactic Database [^3] This is because the observed brightness temperature $\propto DT^{\prime}_{b}/(1+z)$ was assumed by Homan et al. (2021) to correspond to $z=0$ for sources with unknown redshifts, hence the lower limit on $D$ was reported.. The obtained blazar list is denoted as H21+.
@@ 2.2.2 CGRaBS-based source list
Homan et al. (2021) obtained Doppler factors under the assumption that the apparent brightness temperature at $15$ GHz depends only on the blazar jet Doppler factor $D$ and the intrinsic (jet-frame) brightness temperature, which for all blazars in the sample is equal to $4\times 10^{10}$ K, as found according to Kovalev and others (2005). This approach can give only approximate values of the Doppler factor because it does not allow for intrinsic source-to-source variations and, besides, may result in incorrect estimations of $D$ when radio images of the jet and the core of the blazar are superimposed (Kovalev et al., 2025).
Thus, it is of interest to search for blazar-neutrino associations using a sample of sources with their Doppler factors obtained with a different method. Recently, Rodrigues et al. (2024b) modelled time-averaged spectral-energy distributions (SEDs) for a sample of 324 blazars, obtaining a fit to the multi-wavelength SEDs with a leptonic and a leptohadronic model, determining values of the Doppler factor $D$ (among other parameters) for each source. The source sample of Rodrigues et al. (2024b) is based on the flux-limited ($F_{8.4\mathrm{\,GHz}}>65$ mJy) CGRaBS catalogue of radio-loud AGNs (Healey et al., 2008) which was cross-matched by Paliya et al. (2017) with sources from various Fermi-LAT catalogues.
We cross-match sources from the list of 324 blazars (Rodrigues et al., 2024b) [^4] Their Doppler factors and redshifts available at https://github.com/xrod/lephad-blazars/blob/main/model_parameters.csv. with the Fermi-LAT LCR (Abdollahi and others, 2023). This results in $N_{b}=295$ sources with known $D$, $z$ and light curves in the LCR. Among them, according to the Fermi-LAT classification (Abdollahi and others, 2020), there are $218$ FSRQs, $72$ BLLs, $2$ BCUs, and $3$ NLSY1s. We denote the compiled source list as R24+.